OpenStudy (anonymous):

Solve. -20= 5/6 x x = please help

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hello can u help

OpenStudy (misty1212):

\[-20=\frac{5}{6}x\\ -20\times \frac{6}{5}=x\] cancel first, multiply last

OpenStudy (anonymous):

thank you

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]